Sunland, Inc. manufactures ergonomically designed computer furniture. Sunland uses a job order costing system. On November 30, the Work in Process Inventory consisted of the following jobs:
Job No. Item Units Accumulated Cost
CC723 Computer caddy 24,000 $900,000
CH291 Chair 15,000 435,000
PS812 Printer stand 30,000 285,000
$1,620,000

On November 30, Sunland’s Raw Materials Inventory account totaled $680,000, and its Finished Goods Inventory totaled $3,456,400. Sunland applies manufacturing overhead on the basis of machine hours. The company’s manufacturing overhead budget for the year totaled $5,000,000, and the company planned to use 1,000,000 machine hours during the year. Through the first eleven months of the year, the company used a total of 870,000 machine hours, total manufacturing overhead amounted to $4,284,500, and Cost of Goods Sold was $8,790,250. For the pursposes of this problem, ignore year-end disposition of over / under applied overhead.

Sunland purchased $648,000 in raw materials in December and incurred the following costs for jobs in process that month:
Job No. Materials
Issued
Machine
Hours
Direct Labor
Hours
Direct Labor
Cost
CC723 $200,000 19,000 12,600 $125,400
CH291 $16,200 4,800 3,600 $48,200
PS812 $218,000 25,000 15,300 $209,500
DS444 $254,000 20,000 20,000 $149,000

The following jobs were completed in December and transferred to the Finished Goods Inventory:
Job No. Item Units
CC723 Computer caddy 24,000
CH291 Chair 15,000
DS444 Desk 8,000
(a) Calculate the total cost of each of the four jobs worked on in December.
CC723 CH291 PS812 DS444
Total Cost $

$

$

$

(b) Calculate the total manufacturing cost for December.
Total manufacturing cost $

(c) Calculate the cost of goods manufactured for December.
Cost of goods manufactured $

(d) Calculate the balance in the Work in Process Inventory account on December 31.
Balance in work in process $

(e) Assume that Sunland sold 15,000 computer caddies, 12,000 chairs, and 4,500 desks in December. Calculate Cost of Goods Sold for the month of December and the ending
Finished Goods Inventory balance on December 31. (For computation purpose round per unit rates to 2 decimal places, e.g. 3.54. Round final answers to 0 decmial places, e.g. 45,000.)
Cost of goods sold $

Finished goods inventory $
Answer #1

Estimated overhead rate =Estimated overhead /estimated machine hours

                                  = 5000000 / 1000000

                                 = $ 5per machine hours

CC723 CH291 PS812 DS444
Direct material 200000 16200 218000 254000
Direct labor cost 125400 48200 209500 149000
Overhead applied ( actual machine hours*overhead rate) 19000*5=95000 4800*5= 24000 25000*5= 125000 20000*5=100000
Total cost incurred in December 420400 88400 552500 503000
Beginning work in process 900000 435000 285000 0
Total cost of Job 1320400 523400 837500 503000

b)

Total manufacturing cost incurred in December = 420400+88400+552500+503000= $ 1564300

c)

Beginning work in process inventory 1620000
Add:Total manufacturing cost incurred in December 1564300
Total work in process 3184300
less:work in process at end December 31 (837500)
cost of goods manufactured 2346800

Cost of Job PS812 (still in process at end) =Beginning cost+ cost during Dec

                        = 285000+ 552500

                        = 837500

d)Work in process inventory December 31: Cost of Job PS812 (still in process at end) =Beginning cost+ cost during Dec

                        = 285000+ 552500

                        = 837500

e)

CC723 CH291 DS444
Total cost of Job[A] 1320400 523400 503000
Number of units 24000 15000` 8000
Unit cost 1320400/24000 =$ 55.02 523400/15000=34.89 62.88
cost of goods sold (unit cost*units sold)[B]

15000*55.02

=825300

12000*34.89

418680

4500*62.88=282960
Finished goods inventory [A-B]

1320400-825300

=495100